If A= 3i^+4j^ and B=7i^+24j^ the vector having same magnitude as B &, parallel A will be

Grade:12th pass

1 Answers

Aksh
6 years ago
A=3i+4j B=7i+24jLet C be the required vectorThereforeC=|B|(unit vector A) (this satisfies the given condition)C=√(7×7+24×24)(A^)C=√(625)(A/|A|)C=25(3i+4j)/√(3×3+4×4)C=25(3i+4j)/5C=25/5(3i+4j)C=5(3i+4j)C=15i+20j